OpenStudy (anonymous):

Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x > –2 y ≤ 2x + 7 A. (1, 11) B. (–1, 6) C. (–1, 4) D. (–3, –1)

OpenStudy (anonymous):

ok so what now i have no idea what to do

OpenStudy (jdoe0001):

graph them both shade them both where their shades intersect, any point on that area, is a solution \(\bf x>-2y\le 2x+7\quad \textit{dividing all sides by -2}\\ \quad \\ \cfrac{x}{-2}>\cfrac{-2y}{-2}\le \cfrac{2x+7}{-2}\\ \quad \\ \textit{recall when multiplying/dividing/exponentializing by a negative value}\\ \quad \\ \textit{we have to }{\color{red}{ flip}}\textit{ the inequality sign}\\ \quad \\ \cfrac{x}{-2}>\cfrac{-2y}{-2}\le \cfrac{2x+7}{-2}\implies \cfrac{x}{-2}<\cfrac{-2y}{-2}\ge \cfrac{2x+7}{-2}\\ \quad \\\implies \begin{cases} \cfrac{x}{-2}<y\implies {\color{blue}{ y> \cfrac{x}{-2}}}\\ \quad \\ \bf {\color{blue}{ y\ge \cfrac{2x+7}{-2}}} \end{cases}\)

OpenStudy (anonymous):

ok i think it is a is that right

OpenStudy (jdoe0001):

hmm dunno check your graphs shading

OpenStudy (anonymous):

Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x + y ≥ 5 x – 2y > 8

OpenStudy (anonymous):

A. (6, 2) B. (6, –2) C. (8, –1) D. (6, 1)